# A Privacy-Centric Approach: Scalable and Secure Federated Learning Enabled by Hybrid Homomorphic Encryption

### v1

@@ 3.3 Hybrid Homomorphic Encryption
  • $\mathbf{HHE.KeyGen}$: The key generation algorithm takes as input a security parameter $\lambda$ and outputs a HE public/private key pair ($\mathsf{pk}$/$\mathsf{sk}$) and a HE evaluation key ($\mathsf{evk}$).
@@ 5 Methodology
[…]
@@
[…]
@@ 6 Experiment
  • Neural Network: We use a fully connected neural network that consists of two linear layers (without bias terms) and a ReLU activation function in between. The network takes a batch of flattened MNIST images, each with length 784, processes it through a hidden layer of 32 neurons, and produces 10 output values (one for each digit 0-9). The plaintext weights are then flattened into 1D vectors before being encrypted in HE and HHE domains. The values in these weight vectors are floating point in the range of $[-1,1]$. The weight vectors are in the JSON format for plaintext version, or in binaries for the encrypted versions, for data persistence and communication.
@@
For each client, each linear layer that is encrypted symmetrically takes 524 KB in disk, making the whole 2 linear layer neural network 1.048 MB in total. The HE-averaged model output by the server after the HHE protocol takes 16.8 MB in size (8.4 MB per encrypted layer). After aggregation, the server will send the HE encrypted model to each client to decrypt. Hence, for each client, the total communication cost for each round is $1.048+16.8=17.848\ \text{MB}$. And for 3 clients, the total communication cost for one FL round under HHE is $17.848*3=53.544\ \text{MB}$. We can see that the majority of communication cost is incurred by the downloading step of the average HE model from the server, while the symmetric model that’s needed to be uploaded by each client is 16 times less heavy in size.
@@ 6.1 Comparison with Plaintext and CKKS-based HE
111155551010101020202020002,00020002{,}0002 , $0004{,}$ $0006{,}$  [^0] Number of ClientsTotal Communication Cost (MB)Communication Cost over 10 RoundsPlainHEHHE

### v2

@@ 3.3 Hybrid Homomorphic Encryption
  • $\mathbf{HHE.KeyGen}$: The key generation algorithm takes as input a security parameter $\lambda$ and outputs a HE public/private key pair ($\mathsf{pk}$$\mathsf{sk}$) and a HE evaluation key ($\mathsf{evk}$).
@@ 5 Methodology
Refer to captionRefer to captionRefer to captionClient 1Client\ 1Client 2Client\ 2Client 3Client\ 3ServerServer MNIST
@@
A1A1A2A2AnAn Plaintext Symmetric Encryption $\widehat{A1}$$\widehat{An}$$\widehat{A2}$$\widehat{A1}$$\widehat{A2}$$\widehat{An}$$\overline{A1}$$\overline{A2}$$\overline{An}$Transciphering HE Aggregation $\widehat{Ai}Ai¯\overline{Ai}$Symmetric encrypted modelHomomorphic encrypted modelPlaintext model modelClientClient Ciphertext
@@ 6 Experiment
  • Neural Network: We use a fully connected neural network that consists of two linear layers (without bias terms) and a ReLU activation function in between. The network takes a batch of flattened MNIST images, each with length 784, processes it through a hidden layer of 32 neurons, and produces 10 output values (one for each digit 0-9). The plaintext weights are then flattened into 1D vectors before being encrypted in HE and HHE domains. The values in these weight vectors are floating point in the range of [1,1][-1,1]. The weight vectors are in the JSON format for plaintext version, or in binaries for the encrypted versions, for data persistence and communication.
@@
For each client, each linear layer that is encrypted symmetrically takes 524 KB in disk, making the whole 2 linear layer neural network 1.048 MB in total. The HE-averaged model output by the server after the HHE protocol takes 16.8 MB in size (8.4 MB per encrypted layer). After aggregation, the server will send the HE encrypted model to each client to decrypt. Hence, for each client, the total communication cost for each round is 1.048+16.8=17.848MB1.048+16.8=17.848\ $\text${MB}. And for 3 clients, the total communication cost for one FL round under HHE is 17.8483=53.544MB17.848*3=53.544\ $\text${MB}. We can see that the majority of communication cost is incurred by the downloading step of the average HE model from the server, while the symmetric model that’s needed to be uploaded by each client is 16 times less heavy in size.
@@ 6.1 Comparison with Plaintext and CKKS-based HE
11551010202002,0002{,}0004,0004{,}0006,0006{,} [^0] Number of ClientsTotal Communication Cost (MB)Communication Cost over 10 RoundsPlainHEHHE